Bonus Problem: (20 points) The composite beam is made of A-36 steel (A) bonded to C83400 red brass (B) and has the cross sectshow all steps no short cuts please
Answer #1

Answer: B4 Brass - 125mm Case A Moment act M: 50 kN.m & Steel 100mm This is composite bar care so we have to change in single→ T- 63.13 X100 + (63.13 x100 x 50+16 12 2100 + (125x100x(50-46. = 33128142.57 +24495086.67 = 1 : 57.623 x 106 mm? Bending stAs = AB Ego in Egna 12 *10*x20 = Gol Hol + zoo ) ~ To = 16.1 MPa] « T5 = 16.1 MPa] Now total stress by moment & Temperature c